Write all the factors of the following numbers: 
20 ??????????????

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

20 = 1 × 20 20 = 2 × 10 20 = 4 ×5 20 = 5 × 4
Stop here since 4 and 5 have occurred earlier
Hence, the factors of 20 are 1, 2, 4, 5 10 and 20

Thanks